# Can World Models Achieve Mapless Path Planning with Semantic Targets?

### v1

@@ 3.2.2 Path Evaluation
We evaluate predicted trajectories on the test split of the Target dataset using a comprehensive suite of metrics that assess accuracy, goal-reaching capability, and path consistency. All metrics assume 2D trajectories represented as sequences of positions, with ground truth $\mathbf{s}^{\mathrm{GT}}=\{s_{1},s_{2},\ldots,s_{T}\}$ and prediction $\hat{\mathbf{s}}=\{\hat{s}_{1},\hat{s}_{2},\ldots,\hat{s}_{T}\}$, where $s_{t},\hat{s}_{t}\in\mathbb{R}^{2}$ and $T$ is the total number of time steps.
@@
where $\sigma=0.6$ m controls the tolerance. $\mathrm{SE}\in[0,1]$ where $1$ indicates perfect alignment.
@@
Evaluates whether the predicted trajectory stays within a progress-dependent corridor around the ground truth path. We uniformly sample $M=20$ reference points along the ground truth trajectory and assign each a variable radius:
@@
where $p_{i}=i/(M-1)$ is the normalized progress, $\sigma_{\min}=0.15$ m, $\sigma_{\max}=0.5$ m, and $\beta=0.25$. A predicted point $\hat{s}_{j}$ is covered if $\min_{i}\|\hat{s}_{j}-s_{i}^{\mathrm{GT}}\|_{2}\leq\sigma_{i}$. The AC score is:
@@
where $N_{\mathrm{p}}$ is the total number of predicted points, $N_{\mathrm{c}}$ is the number covered by the corridor, and $\gamma=5$.
Spatial World Model Open Video Length (s) Metrics WO $\uparrow$
VGGT [wang2025vggtvisualgeometrygrounded] gt_video - 8 - 720p 0.203 0.580 11.08 0.901 0.993 0.783
Sora 2 8 - 720p 1.912 1.289 54.84 0.160 0.788 0.178
Veo 3.1 8 - 720p 3.125 2.432 77.25 0.212 0.450 0.187
Veo 3.1-fast 8 - 720p 2.863 1.798 57.42 0.140 0.691 0.165
Wan2.5-Preview 10 - 720p 2.478 1.596 61.69 0.182 0.873 0.188
Wan2.2-Plus 5 - 1080p 1.377 1.044 45.25 0.240 0.686 0.252
Wan2.2-Flash 5 - 720p 1.362 1.005 38.75 0.292 0.746 0.299
Wan2.1-Plus 5 - 720p 2.782 1.594 53.90 0.090 0.715 0.131
Wan2.1-Turbo 5 - 720p 4.243 1.850 64.21 0.000 1.000 0.046
Wan2.2-5B-base 8 - 720p 3.666 1.636 58.11 0.012 0.944 0.066
Wan2.2-5B-FT§ 8 - 720p 1.320 0.897 31.91 0.261 0.787 0.287
Wan2.2-5B-FT-DA 8 - 720p 1.050 0.816 26.71 0.333 0.810 0.345
Table 1: Evaluation results of Final Displacement Error (FDE), Average Displacement Error (ADE) and Miss Rate (MR), Soft Endpoint (SE), Approach Consistency (AC) and Weighted Overall (WO). Image-to-Video (I2V) models, § fine-tuned, fine-tuned with data augmentation.
@@
with default weights: $w_{\mathrm{ADE}}=0.05$, $w_{\mathrm{FDE}}=0.10$, $w_{\mathrm{MR}}=0.10$, $w_{\mathrm{SE,AC}}=0.65$, and scale parameters $\tau_{\mathrm{ADE}}=\tau_{\mathrm{FDE}}=1.0$ m.
@@ 4.1 Implementation Details
For model fine-tuning we use the LoRA training framework offered by DiffSynth-Studio [DiffSynthStudio_wanvideo], with 8× NVIDIA A800 80 GB GPUs. For inference of the fine-tuned models on the Target Benchmark we utilize one single NVIDIA RTX PRO 6000 Blackwell 96 GB GPU, while closed-source models are accessed directly through their official APIs. All generated videos are produced at 720p or 1080p resolution with durations ranging from 5 to 10 seconds. All evaluation experiments are conducted on a DELL Alienware Aurora R15 workstation equipped with an NVIDIA RTX 4090 GPU.
Figure 6: World model performance comparison with VGGT as world decoder’s spatio-temporal reconstruction tool.
Data World Metrics WO $\uparrow$
Explicit Target gt_video 0.210 0.586 10.71 0.892 0.994 0.778
Sora 2 1.979 1.381 56.93 0.165 0.740 0.177
Veo 3.1 3.275 2.598 80.69 0.172 0.419 0.155
Veo 3.1-f 3.225 1.928 57.37 0.136 0.685 0.162
Wan2.5 2.449 1.656 61.70 0.182 0.826 0.189
Wan2.2-P 1.413 1.028 42.25 0.242 0.724 0.258
Wan2.2-Fl 1.418 1.006 35.69 0.295 0.776 0.305
Implicit Target gt_video 0.193 0.573 11.61 0.913 0.992 0.791
Sora 2 1.819 1.160 51.91 0.154 0.856 0.179
Veo 3.1 2.914 2.200 72.43 0.269 0.493 0.231
Veo 3.1-f 2.355 1.616 57.49 0.144 0.700 0.170
Wan2.5 2.519 1.512 61.67 0.181 0.941 0.188
Wan2.2-P 1.327 1.067 49.46 0.238 0.634 0.242
Wan2.2-Fl 1.283 1.002 43.05 0.288 0.706 0.290
Table 2: Evaluation results with explicit and implicit semantic targets. Wan2.5-I2V-Preview, Image-to-Video (I2V) models.
[…]
gt_video 0.526 0.600 17.19 0.59 0.95 0.558 1.345 0.845 30.22 0.19 0.89 0.249
Sora 2 1.913 1.257 51.33 0.14 0.82 0.177 2.147 1.245 51.04 0.09 0.85 0.142
Veo 3.1 2.607 2.133 78.49 0.20 0.42 0.172 2.916 1.945 55.07 0.16 0.62 0.188
Veo 3.1-fast 3.402 2.150 62.89 0.15 0.59 0.166 2.601 1.715 58.76 0.15 0.62 0.175
Wan2.5-Preview 2.379 1.510 61.93 0.18 0.88 0.187 2.686 1.608 59.52 0.10 0.83 0.132
Wan2.2-Plus 1.409 1.631 73.19 0.25 0.68 0.225 1.621 0.976 36.65 0.18 0.80 0.227
Wan2.2-Flash 1.407 1.572 71.06 0.28 0.71 0.254 1.687 1.006 37.42 0.17 0.81 0.222
Wan2.1-Plus 2.810 2.110 63.08 0.08 0.67 0.118 2.858 1.510 53.70 0.04 0.76 0.102
Wan2.1-Turbo 4.219 1.833 63.45 0.00 1.00 0.047 4.256 1.856 64.35 0.00 1.00 0.046
Wan2.2-TI2V-5B 3.661 1.666 57.05 0.03 0.89 0.085 3.773 1.705 60.27 0.01 0.91 0.066
Wan2.2-5B-FT§ 1.342 0.951 38.53 0.23 0.74 0.259 1.980 1.079 42.76 0.12 0.80 0.176
Wan2.2-5B-FT-DA 1.055 0.875 33.99 0.33 0.75 0.338 1.841 1.023 39.01 0.11 0.83 0.180
Table 3: Evaluation results of SpaTracker and ViPE. Image-to-Video (I2V) models, § Wan2.2-TI2V-5B fine-tuned model, Wan2.2-TI2V-5B model fine-tuned with data augmentation.
@@ 4.2 Evaluation for off-the-shelf Models
Table 1 shows the evaluation results using VGGT as the spatio-temporal reconstruction tool. Among all evaluated off-the-shelf models, Wan2.2-Flash achieves the best overall performance with a weighted overall score of 0.299. Specifically, it obtains the lowest errors in FDE (1.362m), ADE (1.005m), and MR (38.75%), while achieving the highest SE (0.292). Figure 6 visualizes the performance comparison across all models. The ground truth video achieves the highest scores across all metrics, and is still affected by reconstruction errors like the other models.
@@
ADE values range from 1.0m to 2.4m across different models. Veo 3.1 exhibits the highest displacement error (2.432m), while Wan2.2-Flash and Wan2.2-Plus achieve the best accuracy (around 1.0m). For FDE, errors show wider variation, with Wan2.1-Turbo displaying the highest FDE (4.243m).
@@
The Miss Rate varies significantly from 38.75% to 77.25%. Veo 3.1 shows the highest miss rate (77.25%), indicating poor trajectory quality, while Wan2.2-Flash achieves the best (38.75%). Most models score below 0.3 on the Soft Endpoint metric, suggesting challenges in reaching target endpoints accurately.
@@
As shown in Table 2, performance on implicitly defined targets closely matches that on explicit targets, with small, model-specific fluctuations. This indicates that current WMs can understand semantic goals even when they are not explicitly defined.
@@ 4.3 Fine-tuned Models
For WM fine-tuning, we use 325 scenarios from our dataset and sample 121 frames from each video. To expand the dataset size, we apply a shifting-frame augmentation: for each video, we generate four clips by evenly sampling 121 frames starting from different offsets, preserving similar temporal coverage. We choose to fine-tune the open-sourced Wan2.2-TI2V-5B model under two settings: without and with data augmentation. Table 1 evaluates all models on unseen data using VGGT for path reconstruction. The fine-tuned Wan2.2-5B (Wan2.2-5B-FT) improves its score from 0.066 to 0.287. The augmented version (Wan2.2-5B-FT-DA) outperforms the base model by more than 400% and achieves the best overall score.
@@ 4.4 Ablation Study
Fig. 7, Table 1 and Table 3 compare the performance of three spatio-temporal reconstruction tools: VGGT, SpaTracker, and ViPE. VGGT achieves the best results, with the ground truth video obtaining a weighted overall score of 0.783 and the best-performing off-the-shelf world model (Wan2.2-Flash) achieving 0.299. SpaTracker shows slightly lower performance, while ViPE produces the weakest results. The high score achieved by VGGT on ground truth videos (0.783) confirms that decoded trajectories align well with ground truth trajectories, validating our evaluation approach.
Figure 7: Overall score comparison between different spatio-temporal reconstruction tools. Detailed evaluation results with SpaTracker or ViPE can be found in the appendix.
@@
2.2-Flash 8s 1.362 1.005 38.75 0.292 0.746 0.299
6s 1.393 0.762 23.76 0.261 0.579 0.290
4s 1.278 0.713 23.18 0.290 0.574 0.314
2.2-Plus 8s 1.377 1.044 45.25 0.240 0.686 0.252
6s 1.390 0.789 25.36 0.263 0.584 0.291
4s 1.400 0.783 25.91 0.265 0.534 0.292
Table 4: Evaluation results with different path planning horizons, with VGGT. Wan2.2-I2V-Flash, Wan2.2-I2V-Plus.
@@ 4.5 Sensitivity to the Planning Horizon Length
To assess how the path planning horizon influences WM performance, Table 4 reports results for two WMs evaluated at three horizons: 8 s, 6 s, and 4 s. Reducing the horizon from 6 s to 4 s yields an 8% WO improvement for Wan2.2-I2V-Flash, while shortening it from 8 s to 6 s increases the weighted score of Wan2.2-I2V-Plus by more than 15%. Overall, the weighted score consistently improves as the horizon decreases, suggesting that WMs are more reliable when planning on shorter temporal windows.
@@ 5 Discussion
Benchmark Validity and Performance Ceiling.
Ground truth videos achieve a score of 0.783 when decoded with VGGT (Table 1), establishing an upper bound for the WO score. The gap from a perfect score stems from inherent reconstruction limitations: (1) monocular scale recovery introduces systematic errors, and (2) feed-forward pose estimation struggles with motion blur. Critically, this ceiling is not a flaw: it proves that our pipeline correctly distinguishes high-quality inputs (0.783 for ground truth) from poor ones (0.066-0.299 for base models). As reconstruction methods improve, this ceiling can be raised by substituting VGGT with better alternatives, making Target-Bench forward-compatible with future advances.
Current World Model Performance.
The best off-the-shelf WM, Wan2.2-Flash, achieves only 0.299 overall score. This indicates a significant gap between current world model capabilities and reliable path planning. However, qualitative inspection of generated videos reveals that most models correctly understand the semantic target and show plausible motion tendencies. Additionally, our benchmark includes challenging scenarios with implicit semantic targets. Since we evaluate one-time inference without re-planning, WMs show promising potential for improvement.
@@
Target-Bench intrinsically evaluates three capabilities simultaneously: (1) spatio-temporal consistency, (2) semantic reasoning, and (3) geometric path planning accuracy. Blurred frames, temporal discontinuities, or spatial warping will directly result in poor reconstruction and low path accuracy. This holistic assessment fundamentally differs from existing benchmarks that isolate individual dimensions (visual quality, physics consistency) and miss their integration. Our benchmark shows that perceptual realism does not guarantee path planning utility, addressing a critical blind spot in current evaluation practices.
The Power of Targeted Domain Adaptation.
Fine-tuning with merely 325 real-world robot scenarios produces remarkable gains: 423% improvement over the base model, and surpassing all off-the-shelf models. This result challenges conventional wisdom that larger models with more pre-training data always outperform smaller specialized models. Instead, our findings demonstrate that high-quality domain-specific data matter more than scale, for robotic planning tasks. The rapid adaptation (335% gain even without data augmentation) indicates that WMs have strong potential for spatial reasoning and require only targeted exposure to robot-specific tasks, suggesting pathways for deploying WMs in robotics.
@@
Given an observation and semantic goal, WMs predict future frames depicting motion toward the target, and our world-decoder extracts a path for the robot to execute. This has potential for future use in closed-loop robot navigation in unstructured, mapless environments. Recent work [ball2025genie3, yu2025context] shows WMs can explore new spaces while retaining information in latent memory. Combined with our decoder, robots could follow semantic goals using only visual predictions.

### v2

@@ 3.2.2 Path Evaluation
Metrics.
We evaluate the planning results of video world models using a comprehensive suite of metrics that assess accuracy, goal-reaching capability, and path consistency. All metrics assume 2D trajectories represented as sequences of positions, with ground truth $\mathbf{s}^{\mathrm{GT}}=\{s_{1},s_{2},\ldots,s_{T}\}$ and prediction $\hat{\mathbf{s}}=\{\hat{s}_{1},\hat{s}_{2},\ldots,\hat{s}_{T}\}$, where $s_{t},\hat{s}_{t}\in\mathbb{R}^{2}$ and $T$ is the number of time steps.
Figure 6: Illustration of metrics AC and SE with trajectory and endpoint score distribution heatmaps. Plots visualize scores over a fixed set of candidate trajectories. Trajectories with good Approach Consistency can have a bad Endpoint (Far away from the target). Trajectories with a good Endpoint can have a bad Approach Consistency (Not within the oval corridor). Trajectories with a good Overall Score have both good Approach Consistency and Endpoint.
@@
where $\sigma=0.6$ m is the tolerance. $\mathrm{SE}\in[0,1]$ where $1$ indicates perfect alignment.
@@
Evaluates whether the predicted trajectory stays within a progress-dependent corridor around the ground truth path. We uniformly sample $M=20$ reference points along the ground truth trajectory and assign the corridor radius at the $i$-th reference point as $\sigma_{i}$:
@@
where $p_{i}=i/(M-1)$ is the normalized progress, $\sigma_{\min}=0.15$ m and $\sigma_{\max}=0.5$ m are the minimum and maximum corridor radius, and $\beta=0.25$. A predicted point $\hat{s}_{j}$ is covered if $\min_{i}\|\hat{s}_{j}-s_{i}^{\mathrm{GT}}\|_{2}\leq\sigma_{i}$. The AC score is given as:
@@
where $N_{\mathrm{p}}$ is the total number of predicted points, $N_{\mathrm{c}}$ is the number covered by the corridor, and $\gamma=5$. Importantly, this metric design supports non-uniqueness of valid path. The ground truth path is treated as a rough reference rather than a single correct solution, conditioned on the initial heading. As shown in Fig. 6, various trajectories with similar motional tendency can achieve the same level of high score, which means the score is tendency based, instead of trajectory based.
@@
with default weights: $w_{\mathrm{ADE}}=0.075$, $w_{\mathrm{FDE}}=0.125$, $w_{\mathrm{MR}}=0.125$, $w_{\mathrm{SE,AC}}=0.675$, and scale parameters $\tau_{\mathrm{ADE}}=\tau_{\mathrm{FDE}}=1.0$ m. The weight design prioritizes task success; therefore, a dominant weight is assigned to $\mathrm{SE}\!\cdot\!\mathrm{AC}$ to emphasize endpoint proximity and approach consistency, while keeping $\mathrm{ADE}$, $\mathrm{FDE}$, and $\mathrm{MR}$ lightweight to avoid over-emphasizing point-wise trajectory overlapping.
@@ 4 Experiments
Table 3: Evaluation results with VGGT. Image-to-Video (I2V) models, § fine-tuned, fine-tuned with data augmentation.
Spatial World Model Open Video Length (s) Metrics WO $\uparrow$
VGGT [wang2025vggtvisualgeometrygrounded] gt_video - 8 - 720p 0.203 0.580 11.08 0.901 0.993 0.862
Sora 2 8 - 720p 1.912 1.289 54.84 0.160 0.788 0.207
Veo 3.1 8 - 720p 3.125 2.432 77.25 0.212 0.450 0.210
Veo 3.1-fast 8 - 720p 2.863 1.798 57.42 0.140 0.691 0.192
Wan2.5-Preview 10 - 720p 2.478 1.596 61.69 0.182 0.873 0.215
Wan2.2-Plus 5 - 1080p 1.377 1.044 45.25 0.240 0.686 0.290
Wan2.2-Flash 5 - 720p 1.362 1.005 38.75 0.292 0.746 0.341
Wan2.1-Plus 5 - 720p 2.782 1.594 53.90 0.090 0.715 0.157
Wan2.1-Turbo 5 - 720p 4.243 1.850 64.21 0.000 1.000 0.059
Wan2.2-5B-base 8 - 720p 3.666 1.636 58.11 0.012 0.944 0.084
Wan2.2-5B-FT§ 8 - 720p 1.320 0.897 31.91 0.261 0.787 0.330
Wan2.2-5B-FT-DA 8 - 720p 1.050 0.816 26.71 0.333 0.810 0.394
Figure 7: World model performance comparison with VGGT as world-decoder’s spatio-temporal reconstruction tool.
@@ 4.1 Implementation Details
For model fine-tuning we use the LoRA training framework offered by DiffSynth-Studio [DiffSynthStudio_wanvideo], with 8× NVIDIA A800 80 GB GPUs. For inference of the fine-tuned models on the Target Benchmark we utilize one single NVIDIA RTX PRO 6000 Blackwell 96 GB GPU, while closed-source models are accessed directly through their official APIs. All generated videos are produced at 720p or 1080p resolution with durations ranging from 5 to 10 seconds. All evaluation experiments are conducted on a workstation equipped with an NVIDIA RTX 4090 GPU.
@@ 4.2 Evaluation for off-the-shelf Models
Table 3 shows the evaluation results using VGGT as the spatio-temporal reconstruction tool. Among all evaluated off-the-shelf models, Wan2.2-Flash achieves the best overall performance with a weighted overall score of 0.341. Specifically, it obtains the lowest errors in FDE (1.362m), ADE (1.005m), and MR (38.75%), while achieving the highest SE (0.292). Figure 7 visualizes the performance comparison across all models.
@@
ADE values range from 1.0m to 2.4m across different models. Veo 3.1 exhibits the highest displacement error (2.432m), while Wan2.2-Flash and Wan2.2-Plus achieve the best accuracy (around 1.0m). For FDE, errors show wider variation, with Wan2.1-Turbo displaying the highest FDE.
@@
The MR varies significantly from 38.75% to 77.25%. Veo 3.1 shows the highest MR (77.25%), indicating poor trajectory quality, while Wan2.2-Flash achieves the best (38.75%). Most models score below 0.3 on the SE metric, suggesting challenges in reaching target endpoints accurately.
@@
As shown in Table 4, performance on implicitly defined targets closely matches that on explicit targets, with small, model-specific fluctuations. This indicates that current video WMs can understand semantic goals even when they are not explicitly defined.
@@ 4.3 Fine-tuned Models
We use the train split from our dataset for fine-tuning the open-sourced Wan2.2-TI2V-5B model. The text prompt and the first video frame are taken as inputs. We evenly sample 121 frames from each video to form a ground truth frame sequence. The model is fine-tuned under two settings: without and with data augmentation. For data augmentation, we apply a shifting-frame strategy: for each video of a scenario, we generate four clips of 121 frames starting from different offsets, resulting in a four times large training set while preserving similar temporal coverage and reusing the same captions of each sample. Table 3 evaluates all models on unseen data using VGGT for path reconstruction. The fine-tuned Wan2.2-5B (Wan2.2-5B-FT) improves its score from 0.084 to 0.330. The augmented version (Wan2.2-5B-FT-DA) outperforms the base model by more than 469% and achieves the best overall score.
Table 4: Evaluation results with explicit and implicit semantic targets.
Data World Metrics WO $\uparrow$
Explicit Target gt_video 0.210 0.586 10.71 0.892 0.994 0.857
Sora 2 1.979 1.381 56.93 0.165 0.740 0.206
Veo 3.1 3.275 2.598 80.69 0.172 0.419 0.175
Veo 3.1-fast 3.225 1.928 57.37 0.136 0.685 0.188
Wan2.5-I2V-Preview 2.449 1.656 61.70 0.182 0.826 0.216
Wan2.2-I2V-Plus 1.413 1.028 42.25 0.242 0.724 0.298
Wan2.2-I2V-Flash 1.418 1.006 35.69 0.295 0.776 0.349
Implicit Target gt_video 0.193 0.573 11.61 0.913 0.992 0.871
Sora 2 1.819 1.160 51.91 0.154 0.856 0.210
Veo 3.1 2.914 2.200 72.43 0.269 0.493 0.259
Veo 3.1-fast 2.355 1.616 57.49 0.144 0.700 0.198
Wan2.5-I2V-Preview 2.519 1.512 61.67 0.181 0.941 0.214
Wan2.2-I2V-Plus 1.327 1.067 49.46 0.238 0.634 0.278
Wan2.2-I2V-Flash 1.283 1.002 43.05 0.288 0.706 0.331
Table 5: Evaluation results of SpaTracker and ViPE. I2V models, § Wan2.2-TI2V-5B fine-tuned model, Wan2.2-TI2V-5B model fine-tuned with data augmentation.
[…]
gt_video 0.526 0.600 17.19 0.59 0.95 0.623 1.345 0.845 30.22 0.19 0.89 0.291
Sora 2 1.913 1.257 51.33 0.14 0.82 0.208 2.147 1.245 51.04 0.09 0.85 0.170
Veo 3.1 2.607 2.133 78.49 0.20 0.42 0.194 2.916 1.945 55.07 0.16 0.62 0.218
Veo 3.1-fast 3.402 2.150 62.89 0.15 0.59 0.191 2.601 1.715 58.76 0.15 0.62 0.202
Wan2.5-Preview 2.379 1.510 61.93 0.18 0.88 0.213 2.686 1.608 59.52 0.10 0.83 0.156
Wan2.2-Plus 1.409 1.631 73.19 0.25 0.68 0.253 1.621 0.976 36.65 0.18 0.80 0.265
Wan2.2-Flash 1.407 1.572 71.06 0.28 0.71 0.284 1.687 1.006 37.42 0.17 0.81 0.260
Wan2.1-Plus 2.810 2.110 63.08 0.08 0.67 0.141 2.858 1.510 53.70 0.04 0.76 0.125
Wan2.1-Turbo 4.219 1.833 63.45 0.00 1.00 0.061 4.256 1.856 64.35 0.00 1.00 0.059
Wan2.2-TI2V-5B 3.661 1.666 57.05 0.03 0.89 0.104 3.773 1.705 60.27 0.01 0.91 0.083
Wan2.2-5B-FT§ 1.342 0.951 38.53 0.23 0.74 0.300 1.980 1.079 42.76 0.12 0.80 0.208
Wan2.2-5B-FT-DA 1.055 0.875 33.99 0.33 0.75 0.385 1.841 1.023 39.01 0.11 0.83 0.214
@@ 4.4 Ablation Study
Fig. 8, Table 3 and Table 5 compare the performance of three spatio-temporal reconstruction tools: VGGT, SpaTracker, and ViPE. VGGT achieves the best weighted overall score of 0.862 with the ground truth videos. SpaTracker shows lower performance, while ViPE produces the weakest results. The high score achieved by VGGT on ground truth videos confirms that decoded trajectories align well with ground truth trajectories, validating our evaluation approach. Notably, this component is replaceable in our framework as temporal reconstruction methods continue to advance.
Figure 8: Overall score comparison between different spatio-temporal reconstruction tools. Detailed evaluation results can be found in the appendix.
Table 6: Evaluation results with different path planning horizons, with VGGT.
@@
Wan2.2-I2V-Flash 8s 1.362 1.005 38.75 0.292 0.746 0.341
6s 1.393 0.762 23.76 0.261 0.579 0.337
4s 1.278 0.713 23.18 0.290 0.574 0.363
Wan2.2-I2V-Plus 8s 1.377 1.044 45.25 0.240 0.686 0.290
6s 1.390 0.789 25.36 0.263 0.584 0.338
4s 1.400 0.783 25.91 0.265 0.534 0.339
@@ 4.5 Sensitivity to the Planning Horizon Length
To assess how the path planning horizon influences WM performance, Table 6 reports results for two WMs evaluated at three horizons: 8 s, 6 s, and 4 s. Reducing the horizon from 6 s to 4 s yields an 8% WO improvement for Wan2.2-I2V-Flash, while shortening it from 8 s to 6 s increases the weighted score of Wan2.2-I2V-Plus by more than 17%. Overall, the weighted score consistently improves as the horizon decreases, suggesting that WMs are more reliable when planning on shorter temporal windows.
@@ 5 Discussion
World Models Performance.
The best off-the-shelf video WM, Wan2.2-Flash, achieves only 0.341 overall score. This indicates a significant gap between current world model capabilities and reliable path planning. However, qualitative inspection of generated videos reveals that most models correctly understand the semantic target and show plausible motion tendencies. Additionally, our benchmark includes scenarios with implicit semantic targets, which makes our benchmark challenging for video world models to harvest good scores.
@@
Target-Bench intrinsically evaluates three capabilities simultaneously: (1) spatio-temporal consistency, (2) semantic reasoning, and (3) geometric path planning accuracy. Blurred frames, temporal discontinuities, or spatial warping will directly result in poor 3D reconstruction which causes low path accuracy. The holistic assessment of our work fundamentally differs from existing benchmarks that isolate individual dimensions (visual quality, physics consistency) and miss their integration.
@@
Given an observation and semantic goal, WMs predict future frames depicting motion toward the target, and our world-decoder extracts a path for the robot to execute. This has potential for future use in closed-loop robot navigation in unstructured, mapless environments. Recent work [ball2025genie3, yu2025context] shows WMs can explore new spaces while retaining information in latent memory. Combined with our decoder, robots could follow semantic goals using only visual predictions. As shown in the appendix, we also show some preliminary results in terms of deployment of video world model with world-decoder on a quadruped robot regarding real-world navigation tasks.